• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dhamilton

    Log entry time 16:03:09 on November16,2008

    Entry number 249235

    keyword=Day Shift Summary

    Day Shift Summary
    David Hamilton (SL)
    Zhihnog Ye (TO)
    Franco Garibaldi (3rd)

    08:00 - Shift starts with MCC struggling to get a stable 8 microamp beam on target during a RICH pedestal run.
    09:05 - Stop run 3555 to change BB HV and prescales for stable 8 microamp running. Started run 3556.
    10:05 - MCC called to inform us that they have a problem with a cavity and beam will be down for around 15 minutes. Data-taking has been stopped.
    10:15 - Acc. problem fixed ... data taking resumed after rebooting run control due ot it freezing up.
    13:10 - Problem with target GUI. Stopped data-taking for 10 minutes followed by 10 minute test run (3563).
    14:10 - increased beam current to 10 and then 12 microamps, whilst watching the dark current on the BB chambers with a view to changing HV settings.
    14:30 - MCC called to say they are having some problems with the RF system
    15:10 - MCC want to send tuned beam - move to BeO target.
    15:50 - After a while with MCC struggling to get any kind of stable beam in the hall, positions now look good on the BeO. Have requested a harp scan on 3A and 3B - both are less than 300 microns.
    16:00 - New shift begins moving the 3He target into place.

    Run List:
    3555 0.8M - unstable beam current
    3556 1.0M
    3557 0.2M - stopped early due to Acc. problem
    3558 0.7M - stopped early to fix HRSL VDC timing
    3560 1.0M
    3561 0.4M - stopped early to investigate trigger timing
    3562 0.4M
    3563 0.8M
    3565 1.0M